# Images are Achilles’ Heel of Alignment: Exploiting Visual Vulnerabilities for Jailbreaking Multimodal Large Language Models

### v1

@@ Abstract
In this paper, we study the harmlessness alignment problem of multimodal large language models (MLLMs). We conduct a systematic empirical analysis of the harmlessness performance of representative MLLMs and reveal that the image input poses the alignment vulnerability of MLLMs. Inspired by this, we propose a novel jailbreak method named HADES, which hides and amplifies the harmfulness of the malicious intent within the text input, using meticulously crafted images. Experimental results show that HADES can effectively jailbreak existing MLLMs, which achieves an average Attack Success Rate (ASR) of 90.26% for LLaVA-1.5 and 71.60% for Gemini Pro Vision. Our code and data will be publicly released.
@@ 1 Introduction
Recently, by leveraging the powerful capacity of large language models (LLMs) [37], a variety of multimodal large language models (MLLMs) [36] have ${}^{{\ddagger}}$, which can process both textual and visual information similarly as that LLMs process textual input. MLLMs have not only shown superior performance in various visual-language tasks but also possess the capability to engage in image-related dialogues with human users [18, 39]. However, MLLMs also confront similar harmlessness challenges that afflict their backbone LLMs.
@@ 3.1 Hiding Harmfulness from Text to Image
Specifically, we replace the harmful keyword or key phrase from each text instruction $t$ with a text-to-image pointer and utilize images to represent it. On the text side, we categorize all the keywords into three classes: objects, concepts, or behaviors. For the keywords falling under the first two categories, the text-to-image pointer is “the object/concept in the image”, while for the keywords denoting behaviors, the text-to-image pointer is “conduct the behavior on”. On the image side, as the keywords may represent abstract concepts or behaviors that are difficult for models to grasp when depicted by real-world images, we employ typography to represent these keywords. As a result, the generation process of MLLMs can be formally given as:
@@ 4.2 Experiment Results
The incorporation of $i_{\text{opt}}$ under the $+$Opt image setting significantly increases the ASR of all models, even more than 30 percent (e.g., LLaVA ($+$33.86%) and Gemini ProV ($+$37.20%)). These results further verify our previous empirical finding: harmful images tend to elicit harmful responses. Moreover, $i_{\text{opt}}$ crafted in Sec. 3.2 helps mitigate the misunderstanding issues observed with LLaVA-1.5 and LLaVA-1.5L. The ASR of these two models increases notably compared to the $+$Text-to-image pointer setting. We attribute such results to that $i_{\text{opt}}$ always describes scenarios relevant to the instruction contents, which provides extra context that helps MLLMs to accurately understand the instruction.
@@ 5 Related Work
Similar to LLMs, MLLMs also suffer from malicious prompts and likely generate harmful content [24, 37]. To explore the harmfulness of MLLMs, some studies try to prob the potential harmfulness of MLLMs with benchmarks, while others employ white- and black-box attacks on MLLMs with adversarial images and prompts. However, existing benchmarks mainly focus on the effectiveness capacities of MLLMs, lacking sufficient consideration of harmlessness evaluation [10, 15, 20]. To bridge this gap, recent studies have developed a red teaming attack benchmark covering different scenarios to systematically analyze the harmlessness of MLLMs [31, 19, 16]. Similar to LLMs, existing work on jailbreaking MLLMs can be categorized into white- and black-box attacks. For white-box attacks on MLLMs, existing work [24, 26, 4, 27, 9] can be further classified into two categories according to attacking objectives: input images and visual embeddings. For input images, recent studies generate adversarial images with constraints of a harmful response set [24, 26, 9] or utilizing a teacher-forcing optimization approach [4]. For visual embeddings, Shayegani et al. [27] generates adversarial images that look harmless but are similar to the embeddings of harmful images, thereby bypassing harmful content filters. In contrast, recent work in black-box attacks jailbreak the harmlessness alignment of MLLMs by employing techniques such as system prompt attacks [35], transferring harmful information into text-oriented images [12], generating adversarial images with surrogate models [38], and maximum likelihood-based jailbreaking [21]. In our work, we first investigate whether the visual input influences the harmlessness alignment of MLLMs and then incorporate promoting engineering, white- and black-box attacks on MLLMs.
@@ 0.A.1 Experimental Setting
In total, we collect 2286 instructions. We shuffle all these instructions while ensuring that each negative instruction is followed by its positive counterpart. This method aims to teach MLLMs to differentiate between harmful and harmless instructions by contrasting them, thereby learning which instructions should be followed. Subsequently, we adopt these instructions to finetune LLaVA-1.5 using LoRA. The resulting LoRA is named as contrastive harmlessness LoRA.
@@ Appendix 0.F Pseudo Code for Image Harmfulness Optimization
1 Initialize conversation history $h=[p_{sys}]$ Initialize optimized image set $I=\emptyset$ $i_{\text{opt}}^{0}=\mathcal{D}(p_{0})$ $I=I\cup i_{\text{opt}}^{0}$ for $k=0$ to $K-1$ do

### v2

@@ Abstract
In this paper, we study the harmlessness alignment problem of multimodal large language models (MLLMs). We conduct a systematic empirical analysis of the harmlessness performance of representative MLLMs and reveal that the image input poses the alignment vulnerability of MLLMs. Inspired by this, we propose a novel jailbreak method named HADES, which hides and amplifies the harmfulness of the malicious intent within the text input, using meticulously crafted images. Experimental results show that HADES can effectively jailbreak existing MLLMs, which achieves an average Attack Success Rate (ASR) of 90.26% for LLaVA-1.5 and 71.60% for Gemini Pro Vision. Our code and data are available at https://github.com/RUCAIBox/HADES.
@@ 1 Introduction
Recently, by leveraging the powerful capacity of large language models (LLMs) [38], a variety of multimodal large language models (MLLMs) [37] have emerged, which can process both textual and visual information similarly as that LLMs process textual input. MLLMs have not only shown superior performance in various visual-language tasks but also possess the capability to engage in image-related dialogues with human users [18, 40]. However, MLLMs also confront similar harmlessness challenges that afflict their backbone LLMs.
@@ 3.1 Hiding Harmfulness from Text to Image
Specifically, we replace the harmful keyword or key phrase from each text instruction $t$ with a text-to-image pointer and utilize images to represent it. On the text side, we categorize all the keywords into three classes: objects, concepts, or behaviors. For the keywords falling under the first two categories, the text-to-image pointer is “the object/concept in the image”, while for the keywords denoting behaviors, the text-to-image pointer is “conduct the behavior in the image on”. On the image side, as the keywords may represent abstract concepts or behaviors that are difficult for models to grasp when depicted by real-world images, we employ typography to represent these keywords. As a result, the generation process of MLLMs can be formally given as:
@@ 4.2 Experiment Results
The incorporation of $i_{\text{opt}}$ under the $+$Opt image setting significantly increases the ASR of all models, even more than 30 percent (e.g., LLaVA ($+$33.86%) and Gemini ProV ($+$37.20%)). These results further verify our previous empirical finding: harmful images tend to elicit harmful responses. Moreover, $i_{\text{opt}}$ helps mitigate the misunderstanding issues observed with LLaVA-1.5 and LLaVA-1.5L. The ASR of these two models increases notably compared to the $+$Text-to-image pointer setting. We attribute such results to that $i_{\text{opt}}$ always describes scenarios relevant to the instruction contents, which provides extra context that helps MLLMs to accurately understand the instruction.
@@ 5 Related Work
Similar to LLMs, MLLMs also suffer from malicious prompts and likely generate harmful content [24, 38]. To explore the harmfulness of MLLMs, some studies try to prob the potential harmfulness of MLLMs with benchmarks, while others employ white- and black-box attacks on MLLMs with adversarial images and prompts. However, existing benchmarks mainly focus on the effectiveness capacities of MLLMs, lacking sufficient consideration of harmlessness evaluation [10, 15, 20]. To bridge this gap, recent studies have developed a red teaming attack benchmark covering different scenarios to systematically analyze the harmlessness of MLLMs [32, 19, 16]. Similar to LLMs, existing work on jailbreaking MLLMs can be categorized into white- and black-box attacks. For white-box attacks on MLLMs, existing work [24, 26, 4, 27, 9, 30] can be further classified into two categories according to attacking objectives: input images and visual embeddings. For input images, recent studies generate adversarial images with constraints of a harmful response set [24, 26, 9] or utilizing a teacher-forcing optimization approach [4]. For visual embeddings, Shayegani et al. [27] generates adversarial images that look harmless but are similar to the embeddings of harmful images, thereby bypassing harmful content filters. In contrast, recent work in black-box attacks jailbreak the harmlessness alignment of MLLMs by employing techniques such as system prompt attacks [36], transferring harmful information into text-oriented images [12], generating adversarial images with surrogate models [39], and maximum likelihood-based jailbreaking [21]. In our work, we first investigate whether the visual input influences the harmlessness alignment of MLLMs and then incorporate promoting engineering, white- and black-box attacks on MLLMs.
@@ 0.A.1 Experimental Setting
To avoid data leakage issues, we only utilize the first 50% data from HADES to generate instructions and utilize the remaining data for evaluation. We totally collect 2286 instructions. We shuffle all these instructions while ensuring that each negative instruction is followed by its positive counterpart. This method aims to teach MLLMs to differentiate between harmful and harmless instructions by contrasting them, thereby learning which instructions should be followed. Subsequently, we adopt these instructions to finetune LLaVA-1.5 using LoRA. The resulting LoRA is named as contrastive harmlessness LoRA.
@@ Appendix 0.F Pseudo Code for Image Harmfulness Optimization
1 Initialize conversation history $h=[p_{sys}]$
2 Initialize optimized image set $I=\emptyset$
3 $i_{\text{opt}}^{0}=\mathcal{D}(p_{0})$
4 $I=I\cup i_{\text{opt}}^{0}$
5
6for $k=0$ to $K-1$ do